OpenStudy (anonymous):

NEED SERIOUS HELP PLEASE .!!!! Input the equation of the given line in standard form. The line including (3, 1) and (-2, 3)

OpenStudy (anonymous):

eqn of a line passing thru (x1,y1) and (x2,y2) is given by y-y1=((y2-y1)/(x2-x1))*(x-x1)

OpenStudy (anonymous):

that didnt help me any

OpenStudy (anonymous):

take (3,1) as (x1,y1) and (-2,3) as (x2,y2) now plug in the values in the formula i mentioned above

OpenStudy (anonymous):

that formula is confusing me .. I completely suck at math .

OpenStudy (anonymous):

carefully look at the parenthesis ur answer shud b y=(-2/5)x+11/5

OpenStudy (anonymous):

y=mx+c is the standard form of a line m=slope of the line c= y intercpt cut by the line

OpenStudy (anonymous):

i thought the formula was y=mx+b ?

OpenStudy (anonymous):

m=(y2-y1)/(x2-x1) u can take it b,c,d or whatever u like since its a constant

OpenStudy (anonymous):

do u get it now?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

y=-(2/5)x+11/5